< Chapter 2 hmk part 2 Problem 2.31 6 of 10 In a sample of families with 6 children each, the distribution of boys and girls is as shown in the following table: Number of famaes 9 | 6 | 147 204 | 15 Number of boys Part A Calculate the chi-square value to test the hypothesis of a boy-to-girl ratio of 1:1 Express your answer using three decimal places. Request Answer ▼ Part B Are the numbers of boys to girls in these families consistent with the expected 1:1 ratio? O no O yes Request Answer Part C Calculate the chi-square value to test the hypothesis of binominal distribution in six-child families. Express your answer using three decimal places. Request Answer Part D Is the distribution of the numbers of boys and girls in the families consistent with the expectations of binomial probability? O ye
Answer #1

Answer:

Based on the given data:

Number of families 9 63 147 204 151 56 10
Number of girls 0 1 2 3 4 5 6
Number of boys 6 5 4 3 2 1 0

Total number of boys = 9*6 + 63*5 + 147*4 + 204*3 + 151*2 + 56*1 + 10*0 = 1927

Total number of girls = 9*0 + 63*1 + 147*2 + 204*3 + 151*4 + 56*5 + 10*6 = 1913

Total number of child = 1927 + 1913 = 3840

Observed Number (O) Expected Number (1:1) (E) (O-E) (O-E)^2/E
Boys 1927 3840/2 = 1920 -7 0.026
Girls 1913 3840/2 = 1920 7 0.026
Sum (Chi Square) 0.051

Chi square value = 0.051

Degree of freedom = number of categories - 1 = 1

Probability value corresponding to a chi square value of 0.051 with 1 degree of freedom is P-Value = 0.821332. The result is not significant at p < 0.05.

Part A:Calculate the chi-square value to test the hypothesis of a boy-to-girl ratio of 1:1 Express your answer using three decimal places.

  • Chi square value = 0.051

Part B Are the numbers of boys to girls in these families consistent with the expected 1:1 ratio?

  • Yes